# Time Travel in LLMs: Tracing Data Contamination in Large Language Models

### v2

@@ Abstract
Data contamination, i.e., the presence of test data from downstream tasks in the training data of large language models (LLMs), is a potential major issue in measuring LLMs’ real effectiveness on other tasks. We propose a straightforward yet effective method for identifying data contamination within LLMs. At its core, our approach starts by identifying potential contamination at the instance level; using this information, our approach then assesses wider contamination at the partition level. To estimate contamination of individual instances, we employ “guided instruction:” a prompt consisting of the dataset name, partition type, and the random-length initial segment of a reference instance, asking the LLM to complete it. An instance is flagged as contaminated if the LLM’s output either exactly or nearly matches the latter segment of the reference. To understand if an entire partition is contaminated, we propose two ideas. The first idea marks a dataset partition as contaminated if the average overlap score with the reference instances (as measured by ROUGE-L or BLEURT) is statistically significantly better with the completions from guided instruction compared to a “general instruction” that does not include the dataset and partition name. The second idea marks a dataset partition as contaminated if a classifier based on GPT-4 with few-shot in-context learning prompt marks multiple generated completions as exact/near-exact matches of the corresponding reference instances. Our best method achieves an accuracy between 92% and 100% in detecting if an LLM is contaminated with seven datasets, containing train and test/validation partitions, when contrasted with manual evaluation by human experts. Further, our findings indicate that GPT-4 is contaminated with AG News, WNLI, and XSum datasets.
@@ 1 Introduction
To address this issue, we propose an inexpensive and robust approach to detect data contamination for a given dataset partition automatically. Importantly, our approach functions under two realistic assumptions: (a) we lack direct access to the pre-training data of the LLMs, and (b) we have limited computational resources. Intuitively, our method starts by identifying potential contamination in individual instances that are drawn from a small random sample of the corresponding dataset partition (we use samples of 10 instances in this work). Using the information obtains from individual instances, our approach then assesses if an entire dataset partition is contaminated.
@@ 3.1.1 Components to Measure Instance-level Contamination
To gauge instance-level contamination, we utilize two distinct methods: the first leverages BLEURT and ROUGE-L scores, while the second draws on few-shot ICL prompting with GPT-4. While each of these methods requires specific components to be employed, the first two components are shared between both. The third component, the general instruction, is exclusive to the first method. For both methods, we begin our process by steering the LLM towards the (potentially contaminated) dataset partition using guided instruction that integrates the dataset name and the partition of interest. Next, we provide the random-length initial segment of a randomly selected instance and its label if it is available. The LLM is then instructed to complete it. For the first method, we repeat this step using general instruction that omits the dataset and partition name. An example of a guided versus a general instruction is depicted in Figure 1. We detail all the required components below.
@@
(1) Guided Instruction—A Means to Navigate the LLM’s Domain. By employing instruction-tuning on top of causal language modeling (CLM; Vaswani et al. (2017); Radford et al. (2018)), LLMs can be guided by human directives (Wei et al. 2022; Sanh et al. 2022; Chung et al. 2022). This serves as a tool for controlling the LLM’s domain using natural language. Thus, we form guided instruction such that it incorporates the dataset and split name in the input prompt, thereby directing the LLM towards the underlying dataset split. A comprehensive list of all the instructions used in this study for different tasks/datasets can be found in Table 5 in Appendix A.
(2) Unraveling Data History with Causal Language Modeling. Primarily, data contamination occurs during the CLM pre-training phase since it constitutes the largest part of training in LLMs and utilizes web data. Without instruction tuning, an LLM only attempts to complete an input prompt based on data seen during the CLM pre-training phase (Ouyang et al. 2022). Notable models that exhibit this behavior include GPT-2 and GPT-3. We, therefore, employ the next token prediction mechanism to trace data history. In particular, we feed the model the variable-length initial segment of a dataset instance, chosen randomly from a particular split, prompting it to finish the partial instance. For labeled instances, we integrate the corresponding labels in the input prompt. This reflects that if an instance was ingested during the LLM’s pre-training, its label was ingested too.
@@ 3.1.2 Measuring Instance-level Contamination
GPT-4 Evaluation: While both BLEURT and ROUGE-L quantify the overlap between the generated and reference instances, they fall short of pinpointing near-exact matches. To bridge this gap, we adopt few-shot ICL prompting (Brown et al. 2020) to dictate the detection of exact/near-exact matches based on human judgments (see Section  4: Human Evaluation for our definition of a near-exact match). Specifically, this method includes a few representative examples of exact and near-exact matches—sourced from human evaluations—in the prompt, which are used to assess all other generated completions. We chose GPT-4 for this task as it requires no specialized prompting technique (Bubeck et al. 2023), enhancing the reliability of its results. A visual representation of the few-shot ICL prompt used in our study can be seen in Figure 3 in Appendix B. Also, detailed examples, including their ROUGE-L and BLEURT scores, as well as both human and GPT-4 few-shot ICL evaluations, are listed in Table 6 in Appendix C.
@@ 3.2 Detecting Partition-level Contamination
Algorithm 1: A dataset partition is labeled as contaminated if the average overlap score (as provided by BLEURT and ROUGE-L) between the reference instances and generated texts with the guided instruction on a sample of ten instances is statistically significantly better than those produced by general instruction under a non-parametric bootstrap resampling test. [^2] Details of our bootstrap resampling method can be found in Appendix D.
@@ 3.3 Instance Replication: A Valid Approach to Detect Data Contamination
To validate our choice for the hyperparameters used in Algorithm 2, i.e., the number of exact/near-exact matches to declare contamination, we performed a controlled study in which an LLM is contaminated on purpose with several datasets. To this end, we used the GPT-3.5 base model and a subset of the train partition of the following datasets (one dataset from each task in question): AG News, RTE, and XSum. Note that all these partitions were marked as uncontaminated for GPT-3.5 by the human evaluators (see Table 4 and Section 4: Human Evaluation). To mimic the LLM’s pre-training on web data, we retained only minimal metadata about the datasets as they appear on the web when scraped. In particular, we used: the dataset title, the partition name, and the entire instance. [^3] All data formats used for the contamination of GPT-3.5 are detailed in Table 9 in Appendix E. Following training, we evaluate the generated completions by our best-performing technique (Algorithm 2: GPT-4 ICL) (see Table 3). Figure 2 visualizes the generated replicas before and after contamination in one of our experiments when guided instruction is utilized. [^4] Further examples are provided in Table 10 in Appendix F. In addition, Table 2 summarizes our findings from this study. The key conclusion of this experiment is that the contaminated LLM generated at least one exact match in each setting. This underscores that the replication of even one exact match stands as a robust and undeniable indicator of contamination. [^5] Details on the continued training of the GPT-3.5 base model is presented in Appendix E.
As a second experiment, we employed GPT-4 and the GSM8k dataset (Cobbe et al. 2021). This choice was motivated by OpenAI’s technical report on GPT-4, which indicates contamination from its train split (OpenAI 2023). Given that this dataset comprises mathematical problems, our objective is to replicate the questions in the dataset while withholding their corresponding answers. [^6] An example of this replication process is provided in Table 10 in Appendix F. Table 2 reports our results from this experiment. Our results highlight that contamination is not solely identified through exact matches; near-exact matches are also indicative. To account for the probabilistic nature of LLMs, we set a threshold of two for the minimum number of near-exact matches to indicate contamination. As shown, this is supported by the data.
@@ 5 Results and Discussion
(4) Last but not least, the human evaluation reveals that the train and test/validation splits of both the AG News and WNLI datasets were included in GPT-4’s pre-training data. However, for IMDB and RTE, only the training partitions were incorporated, while for XSum, only the test split was leaked. For GPT-3.5, the only data exposure was the test partition of the XSum dataset. These findings confirm that, despite their creators’ efforts, today’s LLMs have ingested NLP datasets. We hope that this observation informs the design of better scientific experiments with LLMs in the NLP space in the future.
Table 4: An assessment of our proposed methods in contrast to ChatGPT-Cheat? method. We evaluate Algorithm 1 using BLEURT and ROUGE-L, as well as Algorithm 2 which relies on GPT-4 decisions via few-shot ICL prompting. The evaluations are performed on 10 instances randomly drawn from each split of a particular dataset, with GPT-4 and GPT-3.5 serving as the LLMs that are investigated. Partition-level contamination is represented in the following ways: (1) While asterisks (*) indicate statistically significant differences between the completions produced by guided and general instructions (as measured by BLEURT and ROUGE-L), underlined numbers indicate settings that align with human evaluations (Algorithm 1). (2) A single tick (✓) points to the presence of at least one exact match, while a double tick (✓✓) signals the identification of two or more near-exact matches (Algorithm 2). A cross sign (×) denotes that neither of the aforementioned conditions were met. For the ChatGPT-Cheat? method, this cross sign indicates that the model’s output does not contain any specific information about the first instances of the dataset partition upon the request to generate them. For the same method, the question mark (?) highlights partitions that are deemed suspicious.
@@ 6 Conclusion
Our evaluation spanned 28 different settings, including seven datasets along with their respective train and test/validation partitions and two LLMs: GPT-4 and GPT-3.5. Our findings indicated that while the replication technique via guided instruction is notably effective, the most accurate evaluation approach that was closely aligned with human judgments for detecting data contamination was the few-shot in-context learning prompt with GPT-4, which integrates a few example instances from human assessments in the input prompt. This method yielded a success rate in pinpointing data contamination across 14/14 scenarios for GPT-4 and 13/14 for GPT-3.5.
@@ Appendix A List of All Guided and General Instructions
Table 5: A comprehensive list of all guided and general instructions used in our experiments. Placeholders include: {split_name} for the partition (or split) name; {dataset_name} for the dataset name; {input} for the first part of the dataset instance cut at the tail randomly or the whole first sentence in NLI-based datasets; and {label} for the corresponding label of the incomplete input instance.
@@
Classification (single-instance) Instruction: You are provided with the first piece of an instance from the {split_name} split of the {dataset_name} dataset. Finish the second piece of the instance as exactly appeared in the dataset. Only rely on the original form of the instance in the dataset to finish the second piece. Label: {label} First Piece: {input} Second Piece: Instruction: Finish the second piece based on the first piece, such that these two pieces become a single instance with the following label. Label: {label} First Piece: {input} Second Piece:
@@
NLI (paired-instance) Instruction: You are provided with Sentence 1 from the {split_name} split of the {dataset_name} dataset. Finish Sentence 2 as appeared in the dataset. Sentence 2 must exactly match the instance in the dataset. Sentence 1: {input} Label: {label} Sentence 2: Instruction: Finish Sentence 2 based on Sentence 1, such that the following label shows the logical relationship between Sentence 1 and Sentence 2. Sentence 1: {input} Label: {label} Sentence 2:
@@
Summarization (single-instance) Instruction: You are provided with the first piece of a summary from the {split_name} split of the {dataset_name} dataset. Finish the second piece of the summary as exactly appeared in the dataset. Only rely on the original form of the summary in the dataset to finish the second piece. First Piece: {input} Second Piece: Instruction: Finish the second piece based on the first piece, such that these two pieces become a single summary. First Piece: {input} Second Piece:
@@
One-sentence Summary (single-instance) Instruction: You are provided with the first piece of a one-sentence summary from the {split_name} split of the {dataset_name} dataset. Finish the second piece of the summary as exactly appeared in the dataset. Only rely on the original form of the summary in the dataset to finish the second piece. First Piece: {input} Second Piece: Instruction: Finish the second piece based on the first piece, such that these two pieces become a single one-sentence summary. First Piece: {input} Second Piece:
Appendix B Few-shot In-context Learning Prompt

### v3

@@ Abstract
Data contamination, i.e., the presence of test data from downstream tasks in the training data of large language models (LLMs), is a potential major issue in measuring LLMs’ real effectiveness on other tasks. We propose a straightforward yet effective method for identifying data contamination within LLMs. At its core, our approach starts by identifying potential contamination at the instance level; using this information, our approach then assesses wider contamination at the partition level. To estimate contamination of individual instances, we employ “guided instruction:” a prompt consisting of the dataset name, partition type, and the random-length initial segment of a reference instance, asking the LLM to complete it. An instance is flagged as contaminated if the LLM’s output either exactly or nearly matches the latter segment of the reference. To understand if an entire partition is contaminated, we propose two ideas. The first idea marks a dataset partition as contaminated if the average overlap score with the reference instances (as measured by ROUGE-L or BLEURT) is statistically significantly better with the completions from guided instruction compared to a “general instruction” that does not include the dataset and partition name. The second idea marks a dataset partition as contaminated if a classifier based on GPT-4 with few-shot in-context learning prompt marks multiple generated completions as exact/near-exact matches of the corresponding reference instances. Our best method achieves an accuracy between 92% and 100% in detecting if an LLM is contaminated with seven datasets, containing train and test/validation partitions, when contrasted with manual evaluation by human experts. Further, our findings indicate that GPT-4 is contaminated with AG News, WNLI, and XSum datasets. [^1] See the paper’s repo at https://github.com/shahriargolchin/time-travel-in-llms.
@@ 1 Introduction
To address this issue, we propose an inexpensive and robust approach to detect data contamination for a given dataset partition automatically. Importantly, our approach functions under two realistic assumptions: (a) we lack direct access to the pre-training data of the LLMs, and (b) we have limited computational resources. Intuitively, our method starts by identifying potential contamination in individual instances that are drawn from a small random sample of the corresponding dataset partition (we use samples of 10 instances in this work). Using the information obtained from individual instances, our approach then assesses if an entire dataset partition is contaminated.
@@ 3.1.1 Components to Measure Instance-level Contamination
To gauge instance-level contamination, we utilize two distinct methods: the first leverages BLEURT and ROUGE-L scores, while the second draws on few-shot ICL prompting with GPT-4. Each method employs particular components; however, the first two—guided instruction and the next token prediction mechanism—are shared. The third component—general instruction—is exclusive to the first method. For both methods, we begin our process by steering the LLM towards the (potentially contaminated) dataset partition using guided instruction that integrates the dataset name, partition of interest, and the random-length initial segment of a randomly selected instance and its label if it is available. The LLM is then instructed to complete it. For the first method, we repeat this step using general instruction that omits the dataset and partition name. An example of a guided versus a general instruction is depicted in Figure 1. We detail all the required components below.
@@
(1) Guided Instruction—A Means to Navigate the LLM’s Output. By employing instruction-tuning on top of causal language modeling (CLM; Vaswani et al. (2017); Radford et al. (2018)), LLMs can be guided by human directives (Wei et al. 2022; Sanh et al. 2022; Chung et al. 2022). This serves as a tool for controlling the LLM’s output using natural language. Thus, we form guided instruction such that it incorporates the dataset and split name in the input prompt, thereby directing the LLM towards the underlying dataset split. A comprehensive list of all the instructions used in this study for different tasks/datasets can be found in Table 5 in Appendix A.
(2) Next Token Prediction—A Means to Unravel Data History. Primarily, data contamination occurs during the CLM pre-training phase since it constitutes the largest part of training in LLMs and utilizes web data. Without instruction tuning, an LLM only attempts to complete an input prompt based on data seen during the CLM pre-training phase (Ouyang et al. 2022). Notable models that exhibit this behavior include GPT-2 and GPT-3. We, therefore, employ the next token prediction mechanism to trace data history. In particular, we feed the model the variable-length initial segment of a dataset instance, chosen randomly from a particular split, prompting it to finish the partial instance. For labeled instances, we integrate the corresponding labels in the input prompt. This reflects that if an instance was ingested during the LLM’s pre-training, its label was ingested too. [^3] Incorporating labels in the input prompt is essential to account for false positives when generating downstream completions. Illustrations of the impact of label integration on downstream completions are provided in Table 6 in Appendix B.
@@ 3.1.2 Measuring Instance-level Contamination
GPT-4 Evaluation: While both BLEURT and ROUGE-L quantify the overlap between the generated and reference instances, they fall short of pinpointing near-exact matches. To bridge this gap, we adopt few-shot ICL prompting (Brown et al. 2020) to dictate the detection of exact/near-exact matches based on human judgments (see Section  4: Human Evaluation for our definition of a near-exact match). Specifically, this method includes a few representative examples of exact and near-exact matches—sourced from human evaluations—in the prompt, which are used to assess all other generated completions. We chose GPT-4 for this task as it requires no specialized prompting technique (Bubeck et al. 2023), enhancing the reliability of its results. A visual representation of the few-shot ICL prompt used in our study can be seen in Figure 3 in Appendix C. Also, detailed examples, including their ROUGE-L and BLEURT scores, as well as both human and GPT-4 few-shot ICL evaluations, are listed in Table 7 in Appendix D.
@@ 3.2 Detecting Partition-level Contamination
Algorithm 1: A dataset partition is labeled as contaminated if the average overlap score (as provided by BLEURT and ROUGE-L) between the reference instances and generated texts with the guided instruction on a sample of ten instances is statistically significantly better than those produced by general instruction under a non-parametric bootstrap resampling test. [^4] Details of our bootstrap resampling method can be found in Appendix E.
@@ 3.3 Instance Replication: A Valid Approach to Detect Data Contamination
To validate our choice for the hyperparameters used in Algorithm 2, i.e., the number of exact/near-exact matches to declare contamination, we performed a controlled study in which an LLM is contaminated on purpose with several datasets. To this end, we used the GPT-3.5 base model and a subset of the train partition of the following datasets (one dataset from each task in question): AG News, RTE, and XSum. Note that all these partitions were marked as uncontaminated for GPT-3.5 by the human evaluators (see Table 4 and Section 4: Human Evaluation). To mimic the LLM’s pre-training on web data, we retained only minimal metadata about the datasets as they appear on the web when scraped. In particular, we used: the dataset title, the partition name, and the entire instance. [^5] All data formats used for the contamination of GPT-3.5 are detailed in Table 10 in Appendix F. Following training, we evaluate the generated completions by our best-performing technique (Algorithm 2: GPT-4 ICL) (see Table 3). Figure 2 visualizes the generated replicas before and after contamination in one of our experiments when guided instruction is utilized. [^6] Further examples are provided in Table 11 in Appendix G. In addition, Table 2 summarizes our findings from this study. The key conclusion of this experiment is that the contaminated LLM generated at least one exact match in each setting. This underscores that the replication of even one exact match stands as a robust and undeniable indicator of contamination. [^7] Details on the continued training of the GPT-3.5 base model are presented in Appendix F.
As a second experiment, we employed GPT-4 and the GSM8k dataset (Cobbe et al. 2021). This choice was motivated by OpenAI’s technical report on GPT-4, which indicates contamination from its train split (OpenAI 2023). Given that this dataset comprises mathematical problems, our objective is to replicate the questions in the dataset while withholding their corresponding answers. [^8] An example of this replication process is provided in Table 11 in Appendix G. Table 2 reports our results from this experiment. Our results highlight that contamination is not solely identified through exact matches; near-exact matches are also indicative. To account for the probabilistic nature of LLMs, we set a threshold of two for the minimum number of near-exact matches to indicate contamination. As shown, this is supported by the data.
@@ 5 Results and Discussion
(4) Last but not least, the human evaluation reveals that the train and test/validation splits of both the AG News and WNLI datasets were included in GPT-4’s pre-training data. However, for IMDB and RTE, only the training partitions were incorporated, while for XSum, only the test split was leaked. For GPT-3.5, the only data exposure was the test partition of the XSum dataset. These findings confirm that, despite their creators’ efforts, today’s LLMs have ingested NLP datasets. We hope that this observation informs the design of better scientific experiments with LLMs in the NLP space.
Table 4: An assessment of our proposed methods in contrast to ChatGPT-Cheat? method. We evaluate Algorithm 1 using BLEURT and ROUGE-L, as well as Algorithm 2 which relies on GPT-4 decisions via few-shot ICL prompting. The evaluations are performed on 10 instances randomly drawn from each split of a particular dataset, with GPT-4 and GPT-3.5 serving as the LLMs that are investigated. Partition-level contamination is represented in the following ways: (1) While asterisks (*) indicate statistically significant differences between the completions produced by guided and general instructions (measured by BLEURT and ROUGE-L), underlined numbers indicate settings that align with human evaluations (Algorithm 1). (2) A single tick (✓) points to the presence of at least one exact match, while a double tick (✓✓) signals the identification of two or more near-exact matches (Algorithm 2). A cross sign (×) denotes that neither of the aforesaid conditions were met. For the ChatGPT-Cheat? method, this cross sign indicates that the model’s output does not contain any specific information about the first instances of the dataset partition upon the request to generate them. For the same method, the question mark (?) highlights partitions that are deemed suspicious.
@@ 6 Conclusion
Our evaluation spanned 28 different settings, including seven datasets along with their respective train and test/validation partitions and two LLMs: GPT-4 and GPT-3.5. Our findings indicated that while the replication technique via guided instruction is notably effective, the most accurate evaluation approach that was closely aligned with human judgments for detecting data contamination was the few-shot in-context learning prompt with GPT-4, which integrates a few example instances from human assessments in the input prompt. This method yielded a success rate in pinpointing data contamination across 14/14 scenarios for GPT-4 and 13/14 for GPT-3.5.101010Limitations. Data contamination can arise from different sources and manifest in various ways, e.g., direct inclusion of dataset instances, metadata contamination, etc. Our best-performing method for detecting contamination (guided instruction with GPT-4 ICL) does not distinguish between different types of contamination, treating both exact and near-exact replicas of dataset instances as indicators of data contamination. Therefore, we encourage future research that can detect contamination, pinpoint its sources, and identify its various forms.
Acknowledgement
We extend our appreciation to Steven Bethard and Eduardo Blanco for their expert guidance and valuable feedback on the early draft of this paper. This work was partially supported by the Defense Advanced Research Projects Agency under the Habitus program. Mihai Surdeanu declares a financial interest in lum.ai. This interest has been properly disclosed to the University of Arizona Institutional Review Committee and is managed in accordance with its conflict of interest policies.
@@ Appendix A List of All Guided and General Instructions
Table 5: A comprehensive list of all guided and general instructions used in our experiments. Placeholders include: {split_name} for the partition (or split) name; {dataset_name} for the dataset name; {first_piece} for the first part of the dataset instance cut at the tail randomly or the whole first sentence in NLI-based datasets; and {label} for the corresponding label of the incomplete input instance.
@@
Classification (single-instance) Instruction: You are provided with the first piece of an instance from the {split_name} split of the {dataset_name} dataset. Finish the second piece of the instance as exactly appeared in the dataset. Only rely on the original form of the instance in the dataset to finish the second piece. Label: {label} First Piece: {first_piece} Second Piece: Instruction: Finish the second piece based on the first piece, such that these two pieces become a single instance with the following label. Label: {label} First Piece: {first_piece} Second Piece:
@@
NLI (paired-instance) Instruction: You are provided with Sentence 1 from the {split_name} split of the {dataset_name} dataset. Finish Sentence 2 as appeared in the dataset. Sentence 2 must exactly match the instance in the dataset. Sentence 1: {first_piece} Label: {label} Sentence 2: Instruction: Finish Sentence 2 based on Sentence 1, such that the following label shows the logical relationship between Sentence 1 and Sentence 2. Sentence 1: {first_piece} Label: {label} Sentence 2:
@@
Summarization (single-instance) Instruction: You are provided with the first piece of a summary from the {split_name} split of the {dataset_name} dataset. Finish the second piece of the summary as exactly appeared in the dataset. Only rely on the original form of the summary in the dataset to finish the second piece. First Piece: {first_piece} Second Piece: Instruction: Finish the second piece based on the first piece, such that these two pieces become a single summary. First Piece: {first_piece} Second Piece:
@@
One-sentence Summary (single-instance) Instruction: You are provided with the first piece of a one-sentence summary from the {split_name} split of the {dataset_name} dataset. Finish the second piece of the summary as exactly appeared in the dataset. Only rely on the original form of the summary in the dataset to finish the second piece. First Piece: {first_piece} Second Piece: Instruction: Finish the second piece based on the first piece, such that these two pieces become a single one-sentence summary. First Piece: {first_piece} Second Piece:
Appendix B Impact of Label Integration on Downstream Completion
To emphasize the impact of incorporating labels into the input prompt on generating downstream completions and to demonstrate their significance in reducing the generation of false positives, Table 6 presents illustrative examples. These examples compare completions produced when a dataset instance is paired with both a correct and an incorrect label in the input prompt. Specifically, when the LLM is prompted with the initial segment of a reference instance and its correct label, the completion aligns exactly with the subsequent part of that reference instance, resulting in an exact match. In contrast, the completion generated under the same initial segment of the reference instance but paired with an incorrect label shows an inexact match.
Table 6: A side-by-side comparison of completions generated by guided instruction when an initial segment of a reference instance is paired with a correct and an incorrect label in the input prompt. In every example, completion under the correct label produces an exact match for the subsequent segment of the reference instance in the input prompt, whereas that with the incorrect label leads to an inexact match. All completions in these examples are generated by GPT-4, as the base model.
Completion under Correct Label (Exact Match) Completion under Incorrect Label (Inexact Match) Instruction: You are provided with Sentence 1 from the validation split of the WNLI dataset. Finish Sentence 2 as appeared in the dataset. Sentence 2 must exactly match the instance in the dataset. Sentence 1: The dog chased the cat, which ran up a tree. It waited at the top. Label: 1 (entailment) Sentence 2: The cat waited at the top. Instruction: You are provided with Sentence 1 from the validation split of the WNLI dataset. Finish Sentence 2 as appeared in the dataset. Sentence 2 must exactly match the instance in the dataset. Sentence 1: The dog chased the cat, which ran up a tree. It waited at the top. Label: 0 (not entailment) Sentence 2: The dog waited at the top. Instruction: You are provided with the first piece of an instance from the train split of the AG News dataset. Finish the second piece of the instance as exactly appeared in the dataset. Only rely on the original form of the instance in the dataset to finish the second piece. Label: 2 (Business) First Piece: Oil and Economy Cloud Stocks’ Outlook (Reuters) Reuters - Soaring crude prices plus worries about the economy Second Piece: and the outlook for earnings are expected to hang over the stock market next week during the depth of the summer doldrums. Instruction: You are provided with the first piece of an instance from the train split of the AG News dataset. Finish the second piece of the instance as exactly appeared in the dataset. Only rely on the original form of the instance in the dataset to finish the second piece. Label: 3 (Sci/Tech) First Piece: Oil and Economy Cloud Stocks’ Outlook (Reuters) Reuters - Soaring crude prices plus worries about the economy Second Piece: and the outlook for the dollar pressured stocks on Friday, leaving the market nearly unchanged.
Appendix C Few-shot In-context Learning Prompt